OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. 1.Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

\[200(1.03)^{2x}=200((1.03)^2)^x=200(1.0609)^x\] for part 2

undeadknight26 (undeadknight26):

Would this equation work? F(x) = 15(.3)^x